# Energy-Efficient Designs for SIM-Based Broadcast MIMO Systems

### v1

@@ 12          $n\leftarrow n+1$
13
14           until $g(\mathbf{U}^{(n+1)})-g(\mathbf{U}^{(n)})>\epsilon_{1,D}$
15          $\lambda_{D}^{(m+1)}\leftarrow g(\mathbf{U}^{(n)})$ according to (18)
16
17          $m\leftarrow m+1$
@@
19          $\mathbf{U}^{(0)}\leftarrow\mathbf{U}^{(n)}$
20
21           until $\lambda_{D}^{(m+1)}-\lambda_{D}^{(m)}>\epsilon_{2,D}$
@@ 29                    $\mathbf{S}_{\text{opt}}=\hat{\mathbf{S}}$ obtained from [25]
$\hat{\mathbf{Y}}$ $=\mathbf{I}+\sum_{k=j+1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{(n)H}\mathbf{U}_{k}^{(n)}\mathbf{H}_{k}$
$=\mathbf{I}+\sum_{k=j+1}^{K}\hat{\mathbf{V}}_{j}^{(n)H}\hat{\mathbf{V}}_{j}^{(n)}\triangleq\hat{\mathbf{Y}}_{j}^{(n)}\mathbb{C}^{N_{t}\times N_{t}}.$
@@
$c^{(n)}$ $=\sum\nolimits_{j=1}^{K}\Bigg{[}\ln\Bigl{|}\mathbf{I}+\hat{\mathbf{V}}_{j}^{(n)}\Bigl{(}\hat{\mathbf{Y}}^{(n)}\Bigr{)}^{-1}\hat{\mathbf{V}}_{j}^{(n)H}\Bigr{|}$
$-\operatorname{Tr}(\hat{\mathbf{V}}_{j}^{(n)}\Bigl{(}\hat{\mathbf{Y}}^{(n)}\Bigr{)}^{-1}\hat{\mathbf{V}}_{j}^{(n)H})-\operatorname{Tr}(\mathbf{A}_{j}^{(n)H})\Bigg{]}$
$=h(\mathbf{U}^{(n)})$
$\quad-\sum\nolimits_{j=1}^{K}\operatorname{Tr}(\hat{\mathbf{V}}_{j}^{(n)}\Bigl{(}\hat{\mathbf{Y}}^{(n)}\Bigr{)}^{-1}\hat{\mathbf{V}}_{j}^{(n)H})-\operatorname{Tr}(\mathbf{A}_{j}^{(n)H})$
@@
$=\Bigl{(}\hat{\mathbf{Y}}_{j}^{(n)}\Bigr{)}^{-1}-\Bigl{(}\hat{\mathbf{Y}}_{j-1}^{(n)}\Bigr{)}^{-1}\in\mathbb{C}^{N_{t}\times N_{t}}$
@@
for $j=1,2,\ldots,K-1$. The above equation indeed suggests a recursive method to compute $\Bigl{(}\hat{\mathbf{Y}}_{j}^{(n)}\Bigr{)}^{-1}$ efficiently. Suppose the inverse of $\hat{\mathbf{Y}}_{j+1}^{(n)}$ is known. Then, we only need to compute the inverse of the matrix $\mathbf{I}+\bar{\mathbf{B}}_{j+1}^{(n)}\hat{\mathbf{V}}_{j+1}^{(n)H}\in\mathbb{C}^{N_{r}\times N_{r}}$, which has complexity of $\mathcal{O}(N_{r}^{3})$, to obtain $\Bigl{(}\hat{\mathbf{Y}}_{j}^{(n)}\Bigr{)}^{-1}$. Thus, starting from $\hat{\mathbf{Y}}_{K}^{(n)}=\mathbf{I}$, we can gradually compute $\Bigl{(}\hat{\mathbf{Y}}_{j}^{(n)}\Bigr{)}^{-1}$ for $j=K-1,K-2,\ldots 1$. In this way, we remark that $\mathbf{A}_{j}^{(n)}$ is also obtained easily.
@@
Now, using the above lower bound of $h(\mathbf{U})$, we consider the following approximate problem
[…]
@@
It is important to note that the above problem is a concave-convex fractional program since $\bar{h}(\mathbf{U};\mathbf{U}^{(n)})$ is concave. Thus, Dinkelbach’s method can be applied to find the optimal solution, which leads to the following parameterized problem
@@
$\underset{\mathbf{U}_{k}}{\operatorname{maximize}}f(\mathbf{U})=c^{(n)}+\sum_{j=1}^{K}2\mathfrak{R}\Bigl{(}\operatorname{Tr}(\mathbf{B}_{j}^{(n)}\mathbf{U}_{j}^{H})\Bigr{)}$ (29)
$-\sum_{j=1}^{K}\operatorname{Tr}\bigl{(}\mathbf{A}_{j}^{(n)H}\sum\nolimits_{k=j}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\bigr{)}$ (30)
$-\lambda_{D}\Bigl{(}\sum_{j=1}^{K}\operatorname{Tr}(\mathbf{U}_{j}^{H}\mathbf{U}_{j})+N_{t}P_{c}+P_{0}+LNP_{s}\Bigr{)}$ (31)
@@
where $\lambda_{D}>0$ is a given parameter. It is important to note that the above optimization problem can be solved independently for each $\mathbf{U}_{j}$, which admits the following closed-form solution
$\mathbf{U}_{j}=\mathbf{B}_{j}^{(n)}\Bigl{(}\mathbf{H}_{j}\sum_{l=1}^{j}\mathbf{A}_{l}^{(n)H}\mathbf{H}_{j}^{H}+\lambda\mathbf{I}\Bigr{)}^{-1},\>j=1,2,\ldots K.$ (32)
@@
Let $\mathbf{S}^{\ast}$ be the optimal solution to (15). Obviously, if $\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k}^{\ast})\leq P_{\max}$, then $\mathbf{S}^{\ast}$ is also optimal to (12a). On the other hand, if $\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k}^{\ast})>P_{\max}$, it is straightforward to see that the optimal solution to (12a) is obtained by solving the sum-rate maximization problem with the sum power constraint, which is defined as:
@@ III-B SIM Phase Shift Optimization
Since the power consumption does not depend on the channel matrix, the phase shift optimization can improve the EE by increasing the sum-rate of the considered system. Therefore, for fixed $\mathbf{S}$, the SIM phase shift optimization problem is formulated as
@@
$\operatorname{subject~{}to}\quad$ $\left|\boldsymbol{\phi}\right|=1.$ (35b)
@@
where $u_{n}$ is the appropriate step size. The gradient of $\kappa(\boldsymbol{\phi})$ w.r.t. the phase shifts of the SIM is determined by the gradients of $\kappa(\boldsymbol{\phi})$ w.r.t. the phase shifts of the constituent SIM layers as
@@
Differentiating $\kappa(\boldsymbol{\phi})$ w.r.t. $\mathbf{\boldsymbol{\Phi}}^{l}$, we obtain
@@
in the previous equation, it can be written as
@@
Since all the elements of $\boldsymbol{\phi}$ have the unit amplitude, the projection $\mathcal{P}_{\phi}(\boldsymbol{\phi})$ is defined by
@@
Finally, the the EE optimization algorithm for a SIM-aided broadcast system with DPC precoding is outlined in Algorithm 2.
@@ 4     Call Algorithm 1 to obtain $\mathbf{S}^{(n+1)}$
10          if $\kappa(\boldsymbol{\phi}^{(n+1)})<\kappa(\boldsymbol{\phi}^{(n)})+\delta||\boldsymbol{\phi}^{(n+1)}-\boldsymbol{\phi}^{(n)}||^{2}$ then
@@ 12               $u_{n}\leftarrow\rho u_{n}$
16                until $\kappa(\boldsymbol{\phi}^{(n+1)})\geq\kappa(\boldsymbol{\phi}^{(n)})+\delta||\boldsymbol{\phi}^{(n+1)}-\boldsymbol{\phi}^{(n)}||^{2}$
@@ 19               $n\leftarrow n+1$
20
21                until convergence of $\eta$ in (12a)
@@
IV Proposed Solution to Linear Precoding
In this section, we propose an optimization method for the EEmax problem with linear precoding. Similarly as in the previous section, the precoding matrices are found by closed-form expressions, which are derived by implementing Dinkelbach’s method, while the optimal phase shifts for the SIM meta-elements are optimized by a gradient-based method.
@@ IV-A Precoding Matrix Optimization
The precoding matrix optimization, for fixed $\boldsymbol{\phi}$, in (13) in fact reduces to the EEmax problem in conventional MIMO systems. It can be observed that the complexity of the direct optimization of $\mathbf{P}$ is proportional to $N_{t}$, which can cause a significant complexity burden for systems even with moderate $N_{t}$, and thus such a direct optimization method is not practically appealing.
To overcome this issue, we consider an equivalent formulation of (13), introduced in [26], that has a smaller dimension and thus requires a lower computational complexity. Denoting $\mathbf{H}=[\mathbf{H}_{1}^{T}\>\mathbf{H}_{2}^{T}\>\cdots\>\mathbf{H}_{K}^{T}]^{T}\in\mathbb{C}^{KN_{r}\times N_{t}}$, (13) can be equivalently written as
@@
$\operatorname{subject~{}to}$ $\sum_{k=1}^{K}\operatorname{Tr}(\bar{\mathbf{H}}\mathbf{X}_{k}\mathbf{X}_{k}^{H})\leq P_{\max},$ (44b)
@@
$\bar{R}_{k}=$ $\ln\Bigg{|}\mathbf{I}+\bar{\mathbf{H}}_{k}\mathbf{X}_{k}\mathbf{X}_{k}^{H}\bar{\mathbf{H}}_{k}^{H}$
×$\times\left(\mathbf{I}+\sum\nolimits_{j=1,j\neq k}^{K}\bar{\mathbf{H}}_{k}\mathbf{X}_{j}\mathbf{X}_{j}^{H}\bar{\mathbf{H}}_{k}^{H}\right)^{-1}\Bigg{|},$ (45)
$\mathbf{X}=[\mathbf{X}_{1}\>\mathbf{X}_{2}\>\cdots\>\mathbf{X}_{K}]\in\mathbb{C}^{KN_{r}\times KN_{r}}$ are new optimization variables, and $\bar{\mathbf{H}}_{k}=\mathbf{H}_{k}\mathbf{H}^{H}\in\mathbb{C}^{N_{r}\times KN_{r}}$ is the $k$-th sub-matrix of $\bar{\mathbf{H}}=\mathbf{H}\mathbf{H}^{H}\in\mathbb{C}^{KN_{r}\times KN_{r}}$. The equivalence between (13) and (44) is a result of [26, Prop. 2], and the optimal solutions of the two problems are related as $\mathbf{P}_{k}=\mathbf{H}^{H}\mathbf{X}_{k}$. We remark that, comparing the size of $\mathbf{X}$ and $\mathbf{P}=[\mathbf{P}_{1},\mathbf{P}_{2},\dots,\mathbf{P}_{K}]\in\mathbb{C}^{N_{t}\times KN_{r}}$, the equivalent formulation in (44) can significantly reduce the number of optimization variables for systems with large $N_{t}$ (i.e., $N_{t}\gg N_{r}$). In the sequel, similar to the development of Algorithm 1, we first drop (44b) and derive a solution for the unconstrained case. The solution for the constrained case follows immediately.
Upon close inspection, we can observe that the denominator of (44) is a quadratic convex function, while the numerator of (44) is neither convex nor concave function. As a result, finding the solution of (44) is not a trivial task. To find an efficient method for solving (44), we again exploit the inequality, given in (22), to obtain a lower bound of the achievable rate in (45). Utilizing the identity $\left|\mathbf{I}+\mathbf{Z}_{k}\mathbf{Z}_{k}^{H}\mathbf{Y}_{k}\right|=\left|\mathbf{I}+\mathbf{Z}_{k}^{H}\mathbf{Y}_{k}\mathbf{Z}_{k}\right|$ to reformulate (45), the lower bound of the achievable rate of user $k$ can be expressed as

### v2

@@ 17    $n\leftarrow n+1$
19     until $\lambda^{(n)}-\lambda^{(n-1)}>\epsilon_{D}$
@@ 27        $\mathbf{S}_{\text{opt}}=\hat{\mathbf{S}}$ obtained from [23]
$\hat{\mathbf{Y}}$ $=\mathbf{I}+\sum\nolimits_{k=j+1}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{(n)H}\mathbf{U}_{k}^{(n)}\mathbf{H}_{k}$
$=\mathbf{I}+\sum\nolimits_{k=j+1}^{K}\hat{\mathbf{V}}_{j}^{(n)H}\hat{\mathbf{V}}_{j}^{(n)}\triangleq\hat{\mathbf{Y}}_{j}^{(n)}\in\mathbb{C}^{N_{t}\times N_{t}}.$
@@
$c^{(n)}=\sum\nolimits_{j=1}^{K}\Bigg[\ln\Bigl|\mathbf{I}+\hat{\mathbf{V}}_{j}^{(n)}\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}\hat{\mathbf{V}}_{j}^{(n)H}\Bigr|$
$-\operatorname{Tr}(\hat{\mathbf{V}}_{j}^{(n)}\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}\hat{\mathbf{V}}_{j}^{(n)H})-\operatorname{Tr}(\mathbf{A}_{j}^{(n)H})\Bigg]$
$=h(\mathbf{U}^{(n)})-\sum\nolimits_{j=1}^{K}\operatorname{Tr}(\hat{\mathbf{V}}_{j}^{(n)}\Bigl(\hat{\mathbf{Y}}^{(n)}\Bigr)^{-1}\hat{\mathbf{V}}_{j}^{(n)H})-\operatorname{Tr}(\mathbf{A}_{j}^{(n)H})$
@@
$=\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}-\Bigl(\hat{\mathbf{Y}}_{j-1}^{(n)}\Bigr)^{-1}\in\mathbb{C}^{N_{t}\times N_{t}}.$
@@
for $j=1,2,\ldots,K-1$. The above equation indeed suggests a recursive method to compute $\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}$ efficiently. Suppose the inverse of $\hat{\mathbf{Y}}_{j+1}^{(n)}$ is known. Then, we only need to compute the inverse of the matrix $\mathbf{I}+\bar{\mathbf{B}}_{j+1}^{(n)}\hat{\mathbf{V}}_{j+1}^{(n)H}\in\mathbb{C}^{N_{r}\times N_{r}}$, which has complexity of $\mathcal{O}(N_{r}^{3})$, to obtain $\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}$. Thus, starting from $\hat{\mathbf{Y}}_{K}^{(n)}=\mathbf{I}$, we can iteratively compute $\Bigl(\hat{\mathbf{Y}}_{j}^{(n)}\Bigr)^{-1}$ for $j=K-1,K-2,\ldots,1$. In this way, $\mathbf{A}_{j}^{(n)}$ is also obtained easily.
@@
Now, using the above lower bound of $h(\mathbf{U})$, we arrive at the following approximate problem
$\text{$\underset{\mathbf{U}_{k}}{\operatorname{maximize}}$ }\frac{\bar{h}(\mathbf{U};\mathbf{U}^{(n)})}{\sum_{k=1}^{K}\operatorname{Tr}(\mathbf{U}_{k}^{H}\mathbf{U}_{k})+N_{t}P_{c}+P_{0}+LNP_{s}}.$ (25)
III-A3 Dinkelbach’s Method: Water-filling Solution
@@
It is important to note that (25) is a concave-convex fractional program since $\bar{h}(\mathbf{U};\mathbf{U}^{(n)})$ is concave. Thus, Dinkelbach’s method can be applied to find the optimal solution, which leads to the following parameterized problem
@@
$\underset{\mathbf{U}_{k}}{\operatorname{maximize}}\;f_{\lambda_{D}}(\mathbf{U})=c^{(n)}+\sum\nolimits_{j=1}^{K}2\mathfrak{R}\Bigl(\operatorname{Tr}(\mathbf{B}_{j}^{(n)}\mathbf{U}_{j}^{H})\Bigr)$
$-\sum\nolimits_{j=1}^{K}\operatorname{Tr}\bigl(\mathbf{A}_{j}^{(n)H}\sum\nolimits_{k=j}^{K}\mathbf{H}_{k}^{H}\mathbf{U}_{k}^{H}\mathbf{U}_{k}\mathbf{H}_{k}\bigr)$
$-\lambda_{D}\Bigl(\sum\nolimits_{j=1}^{K}\operatorname{Tr}(\mathbf{U}_{j}^{H}\mathbf{U}_{j})+N_{t}P_{c}+P_{0}+LNP_{s}\Bigr)$ (26)
@@
where $\lambda_{D}>0$ is a given parameter. Now it is straightforward to see that the above optimization problem can be solved independently for each $\mathbf{U}_{j}$, which admits the following closed-form solution
$\mathbf{U}_{j}=\mathbf{B}_{j}^{(n)}\Bigl(\mathbf{H}_{j}\sum_{l=1}^{j}\mathbf{A}_{l}^{(n)H}\mathbf{H}_{j}^{H}+\lambda_{D}\mathbf{I}\Bigr)^{-1},\>j=1,\ldots K.$ (27)
Substituting (27) into (26) yields
$f_{\lambda_{D}}(\mathbf{U})$ $=\sum_{j=1}^{K}\operatorname{Tr}\Bigl(\mathbf{B}_{j}^{(n)}\Bigl(\mathbf{H}_{j}\sum_{l=1}^{j}\mathbf{A}_{l}^{(n)H}\mathbf{H}_{j}^{H}+\lambda_{D}\mathbf{I}\Bigr)^{-1}\mathbf{B}_{j}^{(n)H}\Bigr)$
$\quad+c^{(n)}-\lambda_{D}\Bigl(N_{t}P_{c}+P_{0}+LNP_{s}\Bigr)$ (28)
Let $\mathbf{H}_{j}\sum_{l=1}^{j}\mathbf{A}_{l}^{(n)H}\mathbf{H}_{j}^{H}=\mathbf{P}_{j}\boldsymbol{\Sigma}_{j}\mathbf{P}_{j}^{H}$ be the eigenvalue decomposition of $\mathbf{H}_{j}\sum_{l=1}^{j}\mathbf{A}_{l}^{(n)H}\mathbf{H}_{j}^{H}$ where $\mathbf{P}_{j}\in\mathbb{C}^{N_{r}\times N_{r}}$ is unitary and $\boldsymbol{\Sigma}_{j}\in\mathbb{C}^{N_{r}\times N_{r}}$ is diagonal. Then, we can rewrite $f_{\lambda_{D}}(\mathbf{U})$ as
$f_{\lambda_{D}}(\mathbf{U})=\sum_{j=1}^{K}\operatorname{Tr}\Bigl(\mathbf{B}_{j}^{(n)}\mathbf{P}_{j}^{H}\Bigl(\boldsymbol{\Sigma}_{j}+\lambda_{D}\mathbf{I}\Bigr)^{-1}\mathbf{P}_{j}\mathbf{B}_{j}^{(n)H}\Bigr)$
$\quad+c^{(n)}+\lambda_{D}\Bigl(N_{t}P_{c}+P_{0}+LNP_{s}\Bigr)$ (29)
$=\sum_{j=1}^{K}\sum_{i=1}^{N_{r}}\frac{d_{ji}}{\sigma_{ji}+\lambda_{D}}+c^{(n)}-\lambda_{D}\Bigl(N_{t}P_{c}+P_{0}+LNP_{s}\Bigr)$ (30)
where $d_{ji}$ is the $i$th diagonal element of $\mathbf{P}_{j}^{H}\mathbf{B}_{j}^{(n)H}\mathbf{B}_{j}^{(n)}\mathbf{P}_{j}$ and $\sigma_{ji}$ is the $i$th diagonal element of $\boldsymbol{\Sigma}_{j}$. Next, we need to find $\lambda_{D}$ such that $f_{\lambda_{D}}(\mathbf{U})=0$, which can be solved using either the bisection method or Newton’s method.
@@
Let $\mathbf{U}_{k}^{\ast}$ be the optimal solution to (25), which is returned by Dinkelbach’s method described above. Then $\mathbf{S}_{k}^{\ast}=(\mathbf{U}_{k}^{\ast})^{{\dagger}}\mathbf{U}_{k}^{\ast}$ is the optimal solution to (14), for $k=1,\ldots,K$. Obviously, if $\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k}^{\ast})\leq P_{\max}$, then $\mathbf{S}^{\ast}=(\mathbf{S}_{1}^{\ast},\mathbf{S}_{2}^{\ast},\dots,\mathbf{S}_{K}^{\ast})$ is also optimal to (11a). On the other hand, if $\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\mathbf{S}_{k}^{\ast})>P_{\max}$, it is straightforward to see that the optimal solution to (11a) is obtained by solving the sum-rate maximization problem with the sum power constraint, which is defined as:
@@ III-B SIM Phase Shift Optimization
Since the total power consumption does not depend on the SIM phase shifts, the SIM phase shift optimization problem for fixed $\mathbf{S}$ is formulated as
@@
$\operatorname{subject~to}\quad$ $\left|\phi_{n}^{l}\right|=1,\forall l,n.$ (33b)
@@
where $\mathcal{P}_{\phi}(\cdot)$ denotes the projection operator that ensures the unit modulus constraint on $\boldsymbol{\phi}$ and $u_{n}$ is the appropriate step size. The gradient of $\kappa(\boldsymbol{\phi})$ w.r.t. the phase shifts of the SIM is determined by the gradients of $\kappa(\boldsymbol{\phi})$ w.r.t. the phase shifts of the constituent SIM layers as
@@
By differentiating $\kappa(\boldsymbol{\phi})$ w.r.t. $\mathbf{\boldsymbol{\Phi}}^{l}$, we obtain
@@
in the previous equation results in
@@
Since all the elements of $\boldsymbol{\phi}$ have unit amplitude, the projection $\mathcal{P}_{\phi}(\boldsymbol{\phi})$ is defined by
@@
[…]
Finally, the EE optimization algorithm for a SIM-aided broadcast system with DPC precoding is given in Algorithm 2.
@@ 4  Call Algorithm 1 to obtain $\mathbf{S}^{(n+1)}$
[…]
@@ 12      $u_{n}\leftarrow\rho u_{n}$
[…]
@@ 21      $n\leftarrow n+1$
22
23       until convergence of $\eta_{\mathrm{dpc}}$ in (11a)
After $\mathbf{S}_{\text{opt}}$ is obtained, optimal covariance matrices in the BC are found using the the BCs-MACs duality.
@@
IV Proposed Solution to LP
In this section, we propose an optimization method for the EEmax problem with LP. Similar to the previous section, the precoding matrices are found in closed-form using Dinkelbach’s method, which also yields a water-filling solution. The phase shifts for the SIM meta-elements are optimized by a gradient-based method.
@@ IV-A Precoding Matrix Optimization
For fixed $\boldsymbol{\phi}$, the precoding matrix optimization in (12) reduces to the EEmax problem in conventional MIMO systems. The structure of our proposed method for precoding matrix optimization as follows:
  • Reduce the dimension of the precoding matrix optimization problem by considering an equivalent reformulation.
  • Apply (19) to obtain a lower bound of the achievable rate for LP.
  • Adopt Dinkelbach’s method to solve the attained problem, which leads to water-filling solution.
IV-A1 Problem Reformulation: Dimension Reduction
It can be observed that the complexity of the direct optimization of $\mathbf{P}$ is proportional to $N_{t}$, which can cause a significant complexity burden even for moderate values of $N_{t}$. Consequently, such a direct optimization method is not practically appealing.
To overcome this issue, we consider an equivalent formulation of (12), introduced in [24], which reduces the problem’s size and computational complexity. Denoting $\mathbf{H}=[\mathbf{H}_{1}^{T}\>\mathbf{H}_{2}^{T}\>\cdots\>\mathbf{H}_{K}^{T}]^{T}\in\mathbb{C}^{KN_{r}\times N_{t}}$, (12) can be equivalently written as
@@ IV-A1 Problem Reformulation: Dimension Reduction
$\operatorname{subject~to}$ $\sum\nolimits_{k=1}^{K}\operatorname{Tr}(\bar{\mathbf{H}}\mathbf{X}_{k}\mathbf{X}_{k}^{H})\leq P_{\max},$ (42b)
@@
$\bar{R}_{k}=\ln\Bigg|\mathbf{I}+\bar{\mathbf{H}}_{k}\mathbf{X}_{k}\mathbf{X}_{k}^{H}\bar{\mathbf{H}}_{k}^{H}\left(\mathbf{I}+\sum\nolimits_{j=1,j\neq k}^{K}\bar{\mathbf{H}}_{k}\mathbf{X}_{j}\mathbf{X}_{j}^{H}\bar{\mathbf{H}}_{k}^{H}\right)^{-1}\Bigg|,$ (43)
$\mathbf{X}=[\mathbf{X}_{1}\>\mathbf{X}_{2}\>\cdots\>\mathbf{X}_{K}]\in\mathbb{C}^{KN_{r}\times KN_{r}}$ are new optimization variables, and $\bar{\mathbf{H}}_{k}=\mathbf{H}_{k}\mathbf{H}^{H}\in\mathbb{C}^{N_{r}\times KN_{r}}$ is the $k$-th sub-matrix of $\bar{\mathbf{H}}=\mathbf{H}\mathbf{H}^{H}\in\mathbb{C}^{KN_{r}\times KN_{r}}$. The equivalence between (12) and (42) is a result of [24, Prop. 2], and the optimal solutions of the two problems are related as $\mathbf{P}_{k}=\mathbf{H}^{H}\mathbf{X}_{k}$. When comparing the size of $\mathbf{X}$ and $\mathbf{P}=[\mathbf{P}_{1},\mathbf{P}_{2},\dots,\mathbf{P}_{K}]\in\mathbb{C}^{N_{t}\times KN_{r}}$, we see that the equivalent formulation in (42) significantly reduces the number of optimization variables, particularly for systems with large $N_{t}$ (i.e., $N_{t}\gg N_{r}$). Following a similar approach as in Algorithm 1, we first derive a solution for the unconstrained case by dropping (42b), with the solution for the constrained case following immediately.
IV-A2 SCA Method: A Tight Lower Bound of Sum Rate
Upon close inspection, we observe that the denominator of (42) is a quadratic convex function, while the numerator of (42) is neither a convex nor concave function. As a result, solving (42) is nontrivial. To develop an efficient solution, we again exploit the inequality, given in (19), to obtain a lower bound of the achievable rate in (43). Using the identity $\left|\mathbf{I}+\mathbf{Z}_{k}\mathbf{Z}_{k}^{H}\mathbf{Y}_{k}\right|=\left|\mathbf{I}+\mathbf{Z}_{k}^{H}\mathbf{Y}_{k}\mathbf{Z}_{k}\right|$ to reformulate (43), the lower bound of the achievable rate of user $k$ can be expressed as